# Multi-scale Cascaded Large-Model for Whole-body ROI Segmentation

### v1

@@ III-D Aggregation in MCFNet
As shown in Fig. 3, the module mainly summarizes the process of multi-scale feature aggregation in the decoding layer stage of two sub-networks, as well as the application of our newly designed adaptive loss aggregation strategy. Multi-scale feature aggregation helps the model utilize both global contextual information and local details, assisting the model in capturing finer image details and complex edge structures. This is particularly important for handling complex small structures in medical images, such as blood vessels and tumor boundaries. Moreover, merging features of different scales helps generate richer and more diverse feature representations, enabling the model to capture more comprehensive image information, effectively dealing with noise and deformation in images, and enhancing the segmentation capability for objects of different sizes. This allows the model to perform more stably across different types of medical images.
Figure 4: Flowchart of the Adaptive-MFA Strategy.
Specifically, the predicted maps output from the four layers of the encoding stage of SEB and B are pairwise fed into the corresponding convolutional blocks for channel-wise processing, obtaining feature maps with the desired specified number of channels. The calculation formula for this part is shown below:
@@
where $E$ refers to the input feature map, $O$ refers to the output feature map, $C$ is the number of input channels, $N$ is the number of label classes, and $K$ is the convolution kernel size. Additionally, $C_{1}=64$, $C_{2}=64$, $C_{3}=128$, $C_{4}=256$. After the convolution block operations, the two sub-networks use bilinear interpolation resizing to match the dimensions of the output feature maps from FCB and SEB, followed by feature map aggregation. Subsequently, the four layers of different scale feature maps are fed into our designed adaptive loss aggregation strategy for prediction map data augmentation, thereby optimizing the training of the entire model. Through the adaptive search optimization accompanying the entire training process, the four layers of feature maps are aggregated to obtain the final output prediction map $Pred$ of the MCFNet model. The aggregation of the four layers of feature maps can be expressed by the following mathematical formula:
@@
where $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ represent the feature maps from the four prediction heads, and $u$, $v$, $w$, $x$ are the corresponding weight coefficients for the four feature maps. In our experiments, we set all the weight coefficients to 1. In the following section, we will provide a detailed introduction to the adaptive loss aggregation strategy.
@@ III-E Adaptive-MFA: Adaptive Multi-scale Feature-Mixing Loss Aggregation
To better optimize model training, we have designed an adaptive loss aggregation strategy. The main idea is to create new prediction maps by combining available prediction maps. By integrating various prediction maps, the model can better handle errors in different situations. Adaptively selecting and weighting different prediction maps can optimize the final prediction results based on performance in different scenarios. This allows the model to better capture complex patterns in the data, thereby improving prediction accuracy.
The entire strategy flowchart is illustrated in the Fig. 4. Specifically, we randomly combine the output prediction maps $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ from the four layers of the decoder stage, resulting in a total of 15 non-empty subset combinations. Subsequently, we divide the 15 combinations into four sets $S_{1}$, $S_{2}$, $S_{3}$ and $S_{4}$, and assign corresponding weights $W_{1}$, $W_{2}$, $W_{3}$ and $W_{4}$ to these sets. These weight coefficients are adaptively adjusted during model training. Regarding the setting of weight coefficients, since $W_{1}+W_{2}+W_{3}+W_{4}=1$ imposes a competitive relationship among the four weight coefficients, it restricts the optimization of model training as it cannot account for all prediction map combinations. In other words, some prediction maps might contribute minimally to the optimization of the training loss but are still crucial for the model training process. Therefore, we discard the $W_{1}+W_{2}+W_{3}+W_{4}=1$ restriction and only stipulate that the initial values satisfy $W_{1}=W_{2}=W_{3}=W_{4}$. The specific contents of the four sets are as follows:
@@ IV Experiments
This section first introduces all 10 datasets used in the experiment, along with the quantitative and qualitative analyses of the experiment. Subsequently, we provide a detailed description of the parameter configurations used in the experiment. Finally, we outline the ablation study conducted for our proposed method.
@@ IV-A Datasets
To validate the efficient segmentation of whole-body risk organs and lesions by MCFNet, we conduct experiments on ten datasets. The detailed descriptions of each dataset are as follows:
@@ IV-A1 HaN (2015)
This is a public dataset for risk organs in the head and neck region, containing CT and MRI images of patients’ head and neck. These patients underwent CT scans and T1-weighted MR imaging to support image-guided radiotherapy planning. The segmentation labels provided by the dataset are stored separately; we preprocessed the segmentation labels of six regions, merging them into one segmentation label file and classifying them in the order of brainstem, mandible, left parotid gland, right parotid gland, left submandibular gland, and right submandibular gland. After preprocessing the dataset, we divide the image data of 17 tumor patients into 14 cases for the training set and 3 cases for the test set, with the training set containing 1924 slices.
@@ IV-A2 SegTHOR (2019)
This dataset is a CT dataset specifically for the segmentation of thoracic risk organs (OARs) and was included in the ISBI 2019 challenge. In this dataset, the risk organs for the segmentation task include the heart, trachea, aorta, and esophagus, which differ in spatial and appearance characteristics. After preprocessing the dataset, we divide the image data of 40 cases into 32 cases for the training set and 8 cases for the test set, with the training set containing 6035 slices.
@@ IV-A3 CHAOS (2019)
As a multi-modal abdominal medical image segmentation dataset, it provides paired multi-modal CT and MRI data along with corresponding annotation information. MRI imaging data is obtained from two different sequences (T1-DUAL and T2-SPIR). Since the CT images only provide liver annotations, and the MRI images provide annotations for four organs (liver, spleen, left kidney, and right kidney), we select MRI imaging of 20 patients with two modalities. After preprocessing the dataset, we divide the image data of 40 cases into 32 cases for the training set and 8 cases for the test set, with the training set containing 1026 slices.
@@ IV-A4 LiTS (2017)
This dataset focuses on the segmentation of the liver and its tumors in CT images. The data were collected from seven different medical centers and were included in the ISBI 2017, MICCAI 2017, and MICCAI 2018 segmentation challenges. From the original dataset of 131 patients, we randomly select data from 40 patients. After preprocessing the images, we obtain 32 cases for the training set and 8 cases for the test set, with the training set containing 4599 slices.
@@ IV-A5 KiTS (2019)
This dataset is a CT dataset for the segmentation of kidneys and their tumors. It was included in the MICCAI 2019 challenge, which aimed to accelerate the development of reliable semantic segmentation methods for kidneys and kidney tumors. All data were sourced from arterial phase abdominal CT scans of 300 kidney cancer patients who underwent partial or radical nephrectomy, with ground truth labels for kidneys and tumors provided. From the 210 publicly available cases, we randomly select data from 80 patients. After preprocessing the images, we obtain 64 cases for the training set and 16 cases for the test set, with the training set containing 4754 slices.
@@ IV-A6 KiTS (2023)
This dataset focuses on the segmentation of kidneys, tumors, and cysts, and is an extension of the KiTS (2019) dataset. It includes additional segmentation labels for cysts and increases the total data volume from 210 cases to 489 cases. This dataset was included in the MICCAI 2023 challenge. From the 489 patients’ image data, we randomly select data from 80 patients for model validation. After preprocessing the data, we divide it into 64 cases for the training set and 16 cases for the test set, with the training set containing 4968 slices.
@@ IV-A7 LUNG (2018)
This dataset is a subtask of the MSD Medical Image Segmentation Decathlon Challenge, aiming to segment lung tumors from CT images, highlighting the challenge of segmenting small targets within a large context. The official dataset includes thin-slice CT scans from 63 non-small cell lung cancer patients. After preprocessing the images, we divide the dataset into 50 cases for the training set and 13 cases for the test set, with the training set containing 1264 slices.
IV-A8 CPCGEA (2023)
This dataset includes MRI data of prostate cancer patients, comprising two modalities: DWI and T2WI sequences. The images suffer from poor clarity and additional noise due to environmental conditions or image sensor issues. Errors during image acquisition can also lead to artifacts such as streaks and shadows. The dataset consists of 172 cases with manually annotated prostate cancer regions. In our experiments, we randomly divide the data into 139 cases for the training set and 33 cases for the test set, with the training set containing 832 slices.
IV-A9 KNEE (2023)
This dataset consists of MRI images of patients’ leg bones, focusing on the segmentation of the femur. It includes MRI image data from 59 patients. After preprocessing the images, we obtain 47 cases for the training set and 12 cases for the test set, with the training set containing 8171 slices.
@@ IV-A9 KNEE (2023)
Figure 5: Training loss trends of the proposed method and seven comparative methods across four datasets.
@@
AutoPET is a large-scale PET/CT dataset focusing on whole-body tumor segmentation, included in the MICCAI 2022 and MICCAI 2023 challenges. The official dataset comprises 1014 paired PET-CT images from 900 patients. Each data instance includes a three-dimensional whole-body FDG-PET image, a three-dimensional whole-body CT image, and a manually annotated three-dimensional tumor mask. Notably, all CT and PET images were acquired in the same scanning session using the same PET/CT scanner, ensuring high anatomical consistency. The dataset primarily covers patients with histologically confirmed malignant melanoma, lymphoma, and lung cancer, as well as negative control patients who underwent FDG-PET/CT scans at the University Hospital of Tübingen and LMU University Hospital in Munich, Germany. In our experiments, we select the PET images and annotations. We randomly divide the data from 80 patients into 64 cases for the training set and 16 cases for the test set, with the training set containing 2558 slices.
@@ IV-B Evaluation Metrics
In all experiments, we comprehensively evaluate the performance of the segmentation models by applying appropriate metrics based on the dataset type.
For segmentation tasks on multi-class datasets, we use two key evaluation metrics: the Dice Similarity Coefficient (DSC) and the 95% Hausdorff Distance (HD95). DSC measures the overlap between the predicted segmentation and the ground truth, with higher values indicating better segmentation performance. HD95 assesses the boundary deviations in the segmentation results, where smaller values reflect higher accuracy in boundary prediction.
For segmentation tasks on single-class datasets, we use four evaluation metrics: DSC, HD95, Recall, and Precision. In addition to DSC and HD95, Recall measures the model’s ability to detect all instances of the target class, showing how well the model captures the entire target region. Precision measures the proportion of correctly identified target regions among all predicted positives, evaluating the model’s ability to avoid false positives. By combining these four metrics, we can more comprehensively and accurately assess the performance of single-class segmentation models.
Figure 6: Visual examples of segmentation on the HaN, SegTHOR, CHAOS, and LiTS datasets.
@@ IV-C Implementation Detail
$I$ is the number of classes, $N$is the total number of voxel. ${Y}_{n,i}$ and ${P}_{n,i}$ are the ground truth and output probability for the $i$-th class at voxel $n$, respectively.
IV-D Comparsion With Other Methods
We employ seven state-of-the-art methods to validate the performance of MCFNet: U-Net, TransUNet, Swin-Unet, UCTransNet, TransCASCADE, Cascaded MERIT, and FSCA-Net. We validate the segmentation performance of our proposed method against seven comparison models across all ten datasets. In Fig. 5, the training loss trends of all methods are illustrated across the four datasets. Fig. 5 shows visual examples of segmentation on all datasets and we highlight incorrectly segmented areas with red boxes. The specific qualitative and quantitative analysis results are as follows:
IV-D1 Experiments on HaN
Our proposed method and seven comparison models segment six regions in this dataset. As shown in the Table I, the Cascaded MERIT model achieves excellent segmentation performance, with an average DSC of 74.71% and an average HD95 of 2.532mm. Our designed MCFNet improves the average DSC by 2.81% and reduces the average HD95 by 0.079mm compared to Cascaded MERIT. The experimental results in Table I show that our model significantly improves the segmentation of the left submandibular gland.
As shown in the first row of Fig. 6, we observe that our proposed method does not exhibit under-segmentation or over-segmentation compared to the comparison models. For example, UNet incorrectly identifies the mandibular region, while Swin-Unet fails to fully recognize the brainstem region, resulting in under-segmentation. Additionally, UCTransNet fails to accurately distinguish between the left and right parotid glands.
@@ IV-D1 Experiments on HaN
IV-D2 Experiments on SegTHOR
Our proposed method and seven comparison models segment the trachea, heart, aorta, and esophagus in this dataset. According to the experimental results in Table II, current state-of-the-art methods such as TransUNet, UCTransNet, and FSCA-Net achieve over 80% in average DSC and reduce the average HD95 to below 5mm. Our proposed method achieves an average DSC of 83.27% and an average HD95 of 4.187mm, with MCFNet obtaining the best performance in segmenting the trachea, heart, aorta, and esophagus.
IV-D3 Experiments on CHAOS
The comparison results are shown in Table III. We can see that the cascade networks TransCASCADE [33] and Cascaded MERIT achieve relatively good segmentation results for the four abdominal organs, with average DSCs of 90.26% and 90.69%, and average HD95s of 3.695mm and 3.914mm, respectively. The traditional TransUNet achieves an average DSC of 91.19%, making it the best-performing model among the comparisons. Our designed model achieves an average DSC of 92.80% and reduces the average HD95 to 3.676mm, significantly outperforming all comparison models.
As shown in the third row of Fig. 6, most comparison models exhibit under-segmentation for the liver region, and Cascaded MERIT also shows ambiguous identification for the spleen and right kidney. Our designed model generally achieves accurate segmentation, but there is still room for improvement in the segmentation of the spleen’s edges.
@@ IV-D3 Experiments on CHAOS
IV-D4 Experiments on LiTS
As shown in the fourth row of Fig. 6, our designed model, like many comparison models, achieves relatively accurate segmentation for the liver and liver tumors. However, there is a gap in the detailed segmentation of the tumor region edges. Additionally, models like UNet, due to the small size of the segmentation target, may fail to recognize the tumor region.
Figure 7: Visual examples of segmentation on the KiTS (2019), KiTS (2023), LUNG, CPCGEA and KNEE datasets.
@@ IV-D4 Experiments on LiTS
IV-D5 Experiments on KiTS (2019)
This dataset focuses on the segmentation of kidneys and kidney tumors. As shown in Table V, our designed model achieves an average DSC that is 2.21% higher than the state-of-the-art model Cascaded MERIT and an average HD95 that is 2.448mm lower. Additionally, MCFNet improves the tumor segmentation DSC by 4.49% compared to TransCASCADE.
As shown in the fifth row of Fig. 7, we demonstrate the segmentation effects of the models on an image containing both the kidney and tumor. Swin-Unet exhibits scattered segmentation in the tumor region, with dispersed pixels. UCTransNet mistakenly identifies some background areas as tumors. Even more problematic, FSCA-Net fails to recognize the tumor region altogether.
@@ IV-D5 Experiments on KiTS (2019)
Figure 8: Visual examples of segmentation on the AutoPET dataset.

### v2

@@ III-D Aggregation in MCFNet
As shown in Fig. 5, the module summarizes the process of multi-scale feature aggregation in the decoding stage and the application of our adaptive loss aggregation strategy. Multi-scale feature aggregation captures global context and fine details, improving segmentation of small, complex structures (e.g., blood vessels and tumors). By merging features from different scales, the model can handle noise, image deformations, and objects of varying sizes, ensuring more stable performance across diverse medical images.
Specifically, the predicted maps output from the four layers of the encoding stage of SEB and FCB are pairwise fed into the corresponding convolutional blocks for channel-wise processing, obtaining feature maps with the desired specified number of channels. The calculation formula for this part is shown below:
@@
where $E$ is the input feature map, $O$ is the output feature map, $C$ is the number of input channels, $N$ is the number of label classes, and $K$ is the convolution kernel size, with $C_{1}=64$, $C_{2}=64$, $C_{3}=128$, and $C_{4}=256$. After convolution, bilinear interpolation is used to resize feature maps from FCB and SEB for aggregation. The four scales of feature maps are then processed by an adaptive loss aggregation strategy for data augmentation, optimizing the training. The final output prediction map $Pred$ is obtained by aggregating these layers through adaptive search optimization during training. The aggregation of the four layers of feature maps can be expressed by the following mathematical formula:
@@
where $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ are the feature maps from the four prediction heads, with $u$, $v$, $w$, $x$ as their corresponding weight coefficients, all set to 1 in our experiments.
Our model produces accurate segmentation maps. Within the aggregation module, combining multi-scale features refines boundaries and improves localization, thereby enhancing overall prediction performance and clinical utility.
Figure 6: Flowchart of the Adaptive-MFA Strategy, which is executed concurrently with the training process.
@@ III-E Adaptive-MFA: Adaptive Multi-scale Feature-Mixing Loss Aggregation
To optimize model training, we design an adaptive loss aggregation strategy, primarily by combining different prediction maps to create new ones. The adaptive selection and weighting of prediction maps help optimize the final results for better prediction accuracy in varying scenarios.
The strategy flowchart in Fig. 6 shows how we combine output prediction maps $P_{1}$, $P_{2}$, $P_{3}$, and $P_{4}$ from the decoder’s four layers, forming 15 non-empty subset combinations. These are divided into four sets $S_{1}$, $S_{2}$, $S_{3}$, with adaptive weights $W_{1}$, $W_{2}$, $W_{3}$, and $W_{4}$ adjusted during training. We remove the constraint $W_{1}+W_{2}+W_{3}+W_{4}=1$ because it hinders prediction maps with minimal contribution from playing an important role in model training. Initially, all weights are set equal: $W_{1}=W_{2}=W_{3}=W_{4}$. The specific contents of the four sets are as follows:
@@ IV Experiments
This section introduces the datasets and implementation settings, then reports detailed quantitative and qualitative results, followed by thorough ablation studies and discussion of the proposed method.
@@ IV-A Datasets
To assess whole-body OARs and lesion segmentation, we evaluate MCFNet on ten datasets spanning multiple anatomical regions and imaging protocols. For each dataset, we report the source, data volume, and annotation scope.
@@ IV-A1 HaN (2015)
This public dataset contains CT and MRI images of patients’ head and neck, used for image-guided radiotherapy planning. The dataset provides segmentation labels for six regions: brainstem, mandible, left and right parotid gland, left and right submandibular gland. After preprocessing, we merged these labels into a single file. The dataset consists of images from 17 tumor patients, with 14 cases used for training (1924 slices) and 3 cases for testing.
@@ IV-A2 SegTHOR (2019)
This CT dataset, part of the ISBI 2019 challenge, is designed for thoracic risk organ segmentation, including the heart, trachea, aorta, and esophagus. These organs differ in spatial and appearance characteristics. After preprocessing, we divided the dataset of 40 cases into 32 for training (6035 slices) and 8 for testing.
@@ IV-A3 CHAOS (2019)
This multimodal abdominal segmentation dataset comprises paired CT–MRI volumes with corresponding annotations. MRI data includes T1-DUAL and T2-SPIR sequences, with annotations for liver, spleen, left kidney, and right kidney. After preprocessing, we split the data of 40 cases into 32 for training (1026 slices) and 8 for testing.
@@ IV-A4 LiTS (2017)
This dataset focuses on liver and tumor segmentation in CT images. It was collected from seven medical centers and included in the ISBI 2017, MICCAI 2017, and MICCAI 2018 challenges. From 131 patients, we randomly select data from 40, with 32 cases for training (4599 slices) and 8 cases for testing.
@@ IV-A5 KiTS (2019)
This dataset focuses on kidney and tumor segmentation in CT images and was part of the MICCAI 2019 challenge. It includes data from 300 kidney cancer patients who underwent nephrectomy. From 210 available cases, we randomly select 80, with 64 cases for training (4754 slices) and 16 cases for testing.
@@ IV-A6 KiTS (2023)
This dataset extends the KiTS (2019) dataset to include cyst segmentation, increasing the total cases from 210 to 489. It was part of the MICCAI 2023 challenge. From 489 cases, we randomly select 80, with 64 cases for training (4968 slices) and 16 cases for testing.
@@ IV-A7 LUNG (2018)
This dataset is part of the MSD Medical Image Segmentation Decathlon Challenge, focusing on lung tumor segmentation from CT scans. It includes thin-slice CT scans from 63 non-small cell lung cancer patients. After preprocessing, we divide it into 50 cases for training (1264 slices) and 13 cases for testing.
@@
Figure 7: Training loss curves of the proposed method and seven comparison methods across four datasets. The stable downward trend of the loss demonstrates the orderly progression of model training.
IV-A8 CPCGEA (2023)
This dataset contains MRI data of prostate cancer patients, including DWI and T2WI sequences. The images are affected by poor clarity and noise, with artifacts like streaks and shadows due to image acquisition issues. It includes 172 cases with annotated prostate cancer regions. We randomly split the data into 139 training cases (832 slices) and 33 test cases.
IV-A9 KNEE (2024)
This dataset consists of MRI images of patients’ leg bones, focusing on femur segmentation. It includes data from 59 patients. After preprocessing, we use 47 cases for training (8171 slices) and 12 cases for testing.
@@ IV-A9 KNEE (2024)
AutoPET is a large-scale PET/CT dataset for whole-body tumor segmentation, included in the MICCAI 2022 and MICCAI 2023 challenges. It contains 1014 paired PET-CT images from 900 patients, focusing on malignant melanoma, lymphoma, and lung cancer. The dataset includes three-dimensional FDG-PET and CT images, with manually annotated tumor masks. We select PET images and annotations from 80 patients, dividing the data into 64 training cases and 16 test cases, with 2558 slices in the training set.
@@ IV-B Evaluation Metrics
We evaluate segmentation models using appropriate metrics based on the dataset type.
For multi-class segmentation, we evaluate using the Dice similarity coefficient (DSC) and the 95th-percentile Hausdorff distance (HD95). DSC measures overlap between the prediction and ground truth (higher is better), while HD95 quantifies boundary error at the 95th percentile (lower is better).
For single-class segmentation, we report DSC, HD95, Recall (Sensitivity), and Precision (Positive Predictive Value). Recall evaluates the model’s ability to capture all target voxels (penalizing false negatives), and Precision measures the proportion of predicted positives that are correct (penalizing false positives). Together, these metrics provide a balanced evaluation of volumetric agreement and boundary accuracy.
Figure 8: Visual examples of segmentation on the HaN, SegTHOR, CHAOS, and LiTS datasets. In the image, different organs or lesions are highlighted using distinct color annotations.
@@ IV-C Implementation Detail
II is the number of classes, NNis the total number of voxel. ${Y}_{n,i}$ and ${P}_{n,i}$ are the ground truth and output probability for the $i$-th class at voxel $n$, respectively.
@@
IV-D Comparsion With Other Methods
We validate the performance of MCFNet across ten datasets using seven state-of-the-art methods: U-Net, TransUNet, Swin-Unet, UCTransNet, TransCASCADE, Cascaded MERIT, and FSCA-Net. Fig. 7 shows the training loss trends of all methods across four datasets, while Fig. 8, 9, 10, and 11 display visual segmentation examples for each dataset, with errors highlighted in red boxes. The specific qualitative and quantitative analysis results are as follows:
IV-D1 Experiments on HaN
Our proposed method and seven comparison models segment six regions in this dataset. As shown in the Table I, the Cascaded MERIT model achieves excellent segmentation performance, with an average DSC of 74.71% and an average HD95 of 2.532mm. Our designed MCFNet improves the average DSC by 2.81% and reduces the average HD95 by 0.079mm compared to Cascaded MERIT. The results in Table I indicate that our model achieves a substantial performance gain on the left submandibular gland.
In the first row of Fig. 8, we see that our method avoids both under-segmentation and over-segmentation. However, UNet misidentifies the mandibular region, Swin-Unet under-segments the brainstem, and UCTransNet struggles to distinguish between the left and right parotid glands.
IV-D2 Experiments on SegTHOR
Our proposed method and seven comparison models segment the trachea, heart, aorta, and esophagus. As shown in Table II, state-of-the-art methods like TransUNet, UCTransNet, and FSCA-Net achieve over 80% DSC and reduce the average HD95 to below 5mm. Our method achieves an average DSC of 83.27% and HD95 of 4.187mm, outperforming others in segmenting these organs.
IV-D3 Experiments on CHAOS
The comparison results are shown in Table III. We can see that the cascade networks TransCASCADE [40] and Cascaded MERIT achieve relatively good segmentation results for the four abdominal organs, with average DSCs of 90.26% and 90.69%, and average HD95s of 3.695mm and 3.914mm, respectively. The traditional TransUNet achieves an average DSC of 91.19%, making it the best-performing model among the comparisons. Achieving 92.80% average DSC and 3.676 mm average HD95, the proposed method surpasses all competing models.
As shown in the third row of Fig. 8, most comparison models under-segment the liver, and Cascaded MERIT also struggles with identifying the spleen and right kidney. Our model achieves generally accurate segmentation, but the spleen’s edges could still be improved.
IV-D4 Experiments on LiTS
@@
As shown in the fourth row of Fig. 8, our model provides accurate segmentation for the liver and tumors. However, fine delineation of tumor boundaries remains challenging. Models such as UNet often struggle to detect small lesions, and even Cascaded MERIT and FSCA-Net exhibit suboptimal performance in capturing hepatic boundary details.
IV-D5 Experiments on KiTS (2019)
The dataset targets segmentation of the kidneys and renal tumors. As shown in Table V, our designed model achieves an average DSC that is 2.21% higher than the state-of-the-art model Cascaded MERIT and an average HD95 that is 2.448mm lower. Additionally, MCFNet improves tumor segmentation DSC by 4.49% over TransCASCADE, and achieves a further 0.21% DSC gain on kidney segmentation compared with Cascaded MERIT.
As shown in the first row of Fig. 9, we show the segmentation effects on an image containing both the kidney and tumor. Swin-Unet has scattered segmentation in the tumor region, UCTransNet mistakenly identifies some background areas as tumors, and FSCA-Net fails to recognize the tumor regions.
@@ IV-D5 Experiments on KiTS (2019)
Figure 9: Visual examples of segmentation on the KiTS (2019), KiTS (2023), LUNG, CPCGEA and KNEE datasets. In the image, different organs or lesions are highlighted using distinct color annotations.